A wire of length 0.050 𝑚 and carries a current 𝐼 is in a horizontal uniform magnetic field B of 2.50 𝑇 which is perpendicular to the rod. The weight of the wire is 0.80 𝑁. What is the value of the current 𝐼 if the magnetic force on the wire equals its weight? 


1
Expert's answer
2021-04-15T07:16:54-0400

"F_A=IBl,"

"F_A=P,\implies"

"I=\frac{P}{Bl},"

"I=\frac{0.8}{2.5\cdot 0.05}=6.4~A."